OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (4, -2) and perpendicular to y = x - 7.

OpenStudy (anonymous):

the equation of any line perpendicular to y = x - 7 is x+y=k

OpenStudy (anonymous):

as it passes through (4,-2) the points will satisfy the equation, 4-2=k

OpenStudy (anonymous):

there is a fraction before x which is 1/3

OpenStudy (anonymous):

so,the equation is x+y=2

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

check your question please..are you sure you have posted it correctly?

OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (4, -2) and perpendicular to y =1/3 x - 7.

OpenStudy (anonymous):

the 3rd one

OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4.

OpenStudy (anonymous):

the next one

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

y=2x+8